OpenStudy (anonymous):

simplify -4(-3y-v)+5(5v-y)

OpenStudy (anonymous):

do you know how to expand those brackets?

OpenStudy (anonymous):

@bluebrandon no

OpenStudy (anonymous):

ok so -4(-3y-v) means you multiply everything inside the brackets by -4 so you would get -4(-3y-v) = (-4)(-3v) - (-4)(v)

OpenStudy (anonymous):

you would do the same thing for the 5(5v-y)

OpenStudy (anonymous):

then all you need to do is collect like terms

OpenStudy (anonymous):

@bluebrandon what's the answer

OpenStudy (anonymous):

what have you got after doing what I said?

OpenStudy (anonymous):

technically use distributive properties and combining like terms

OpenStudy (anonymous):

@bluebrandon is the answer -12v-4v+25vy

OpenStudy (anonymous):

not quite for the first part -4(-3y-v) = (-4)(-3v) - (-4)(v) should give you 12v +4v you shouldn't have any terms with vy either

OpenStudy (anonymous):

i got 7y=29v how is it different please tell me

OpenStudy (anonymous):

12+4v+25y @bluebrandon is that it?

OpenStudy (anonymous):

ok you seem to be having some difficulty I will show you the answer and if you have any questions about what is going on please ask. -4(-3y-v)+5(5v-y) = (-4)(-3y) - (-4)(v) + 5(5v) - 5(y) = 12y + 4v + 25v -5y = 7y +29v

OpenStudy (anonymous):

the last line is where we collected the like terms together

OpenStudy (anonymous):

ok, got it, thx